Fermi National Laboratory
 

Fermilab Shutdown Report for April 14, 2006 - April 21, 2006

 

LINAC

April 14 - 21

  • Finished H.E. LINAC P.M.
  • NTF Patient Treatment (M, W, & Th)
  • LLRF Characterization on LRF3
  • TODAY – Turn off L.E. LINAC
  • Weekend – Power Outage

April 21 - 28

  • MONDAY – Turn on complete LINAC for 400 MeV Beam
  • NTF Patient Treatment (Tu, Th, & F)

Finished Work

  • Cleaned L.E. LINAC Modulators & Rectifiers (LFR1-LRF5)
  • Replace water fittings in Klystron Modulator De-Q’ing circuit
  • Replaced Collimator between L.E. LINAC Tanks 1 & 2
  • L.E. LINAC Water Maintenance
  • Modified Chopper & Moved Quad for future MuCool operations
  • Completed McCool Alignment
  • Annual Pre-Acc Safety Trip Test
  • Cleaned Filters on 8 Ion Pump PS’s
  • 45°/90° Magnet PS Maintenance
  • New I-Anode Bulk Supply On-Line
  • BPM Digitizers Installed (Database work proceeding)

Booster

  • 400 MeV Line Under Vacuum
  • Injection Girder Under Vacuum
  • Power Supplies all connected
  • MI-8 Septa Cables all pulled (24 RG220)
  • Alignment Completed
  • Sump discharge reroute continuing
  • RF repair and maintenance continues

Schedule – Next Week

  • MI-8 Dump on schedule – (tasks remaining)
    • Septa & Kicker PS connections and testing
    • Safety systems checks
  • ORBMP – connect and test PS
  • 400 MeV Power Supply Testing
  • Sump work should finish
  • RF repairs will continue
  • Change Toroid in MI-8 line (break Vacuum!)
  • Power Outages – vacuum issues!

Schedule

  • Startup Planning Meetings (Thursday 13:00 hours – Penthouse)
  • In two weeks – Change Door Cores
  • In two weeks – Search and Secure
  • Try and run ORBMP over weekend – if vacuum is OK

Pbar

On Schedule

Main Injector

  • Continued vacuum repairs in the MI8 line.
  • Continued vacuum repairs at 608 and 620.
  • Continued vacuum repairs at the large aperture quad at 101.
  • Completed the Lambertson bake at 402.
  • Continued installation of large aperture quad at 521,522.
  • Completed testing of the new MI10 kicker loads.
  • Installed Barrier cavity at 604

Recycler

Flying Wires

Pelletron Work

  • New SF6 Skid -Repairs underway
    • Arrived Wednesday
    • Old system has been disconnected and taken out of MI-31
    • Piping work is underway
  • Repair of the following Vacuum Components
    • IPA03 – power supply - Done
    • TPS S05
  • Replacement of shaft Bearings -Repairs underway
  • Repair Collector Current read back- Done
  • Repair Pulser switch -Repairs underway
  • Upgrade Fault protection system -Repairs underway
  • Misc. Maintenance

Critical Items

  • SF6 skid replacement (2-3 weeks)
  • Shafts replacement (1.5 week)
  • Gun modulator repair and testing (2-3 weeks)
  • Venting up gun acceleration tube (1 day) but…
    • Need to limit work under nitrogen purge to 1-2 hours to avoid having to bake
      • Need careful preparation of everyone involved
    • In the event that the tube stays open more than the 1-2 hours limit (unexpected issues like obvious problems in the gun area), we would have to bake à 1 week

Tevatron

Items Finished

  • Smart bolt shimming will finished!
  • Heater on RF cavities moved
  • Vacuum work at F4, E2, B4 finished
  • D1 ready for sign off
  • Separators at A49 and B48

TEL 1

Jobs Not Yet Started

  • Re align low beta quads at B0 week of May 1
  • TEL 1 installation next Monday 4-24
  • HOPs cable pull
  • A3 conning tower replacement, TD will try to do conning towers in A3 if people are available - decision next week

Jobs Not Scheduled But Done

  • Vacuum work A3
  • A38 quad replacement
  • A39 quad roll
  • warm up and vacuum work at D1
  • D16 quad roll
  • F4 dipole and spool replacement
  • TEL 1 removal, repair and replacement
  • Kautzky valve replacement entire ring
  • Flying wire replacement at E17 and E11
  • replace quad stand at B43 and B44
  • replace low beta quad stand at B12
  • move roughing pump at F13
  • unroll 48 magnets around ring
  • measure rolls 13 magnets around ring

SwitchYard

Transfer Hall work continues.

NuMI

  • Horn 2 Repair Completed – Successful ALARA Plan
    • Horn 2 Back in the Chase and T Blocks Installed
    • Horn 2 Water Testing Begins Today
  • R Blocks Over Chase are Off to Allow Target and Horn 1 Inspections
  • Target Chase Chiller Testing Continues at Vendor
    • Still Expecting Return of Chiller by End of Next Week
  • Target Hall Air Dehumidification
    • Desiccant Units Have Arrived at MI-65
    • Job Planning with Prism (Ductwork) Completed – Start Tuesday
    • Scaffolding Erected Below Site Riser 3 (aka EAV 1.5)
    • Electrical Work Continues to Progress
  • Beamline Drain Cleaning Now Considered Substantially Complete
  • Target Profile Monitor Surveyed In After Some Stand Work
  • OTR Drive Problem Fixed. Will Install Next Week.
  • Installation of New Drip Ceiling in Pre-Target Continues

MiniBooNE

  • Horn
    • Supply trips off; experts investigating; will continue testing.
  • RAW System:
    • Week of April 24.
    • May slide depending on horn testing.
    • Require rad tech coverage.
  • Alignment
    • Week of May 1.

Collider Detector at Fermilab

On schedule

D-Zero (D0)

  • Layer 0
    • Complete junction card mounting, cable routing, and associated testing
      • Preliminary test results generally encouraging
        • Capacitance between Layer 0 and beam pipe measured as anticipated
        • Almost all channels (individually) downloaded and readout
          • Noise looks relatively good (so far) on most modules
          • A troublesome module has re-appeared (previously caused trouble at SiDet)
    • Weld stubs on EC beam pipes
  • Level 1 Calorimeter Trigger
    • Complete connecting BLS cables to patch panels and test pulsing BLS cables
    • Begin installation of LVDS cables
    • Continue firmware debugging/development/verification
  • Level 1 Central Track Trigger
    • Complete DFEA2 LVDS output verification
    • Begin connecting/verifying outputs to L1CalTrk and L1Muon
  • Central Fiber Tracker Readout Upgrade
    • AFEII boards released for production

Mechanical

Work Continues

Cryogenics

Cryo Operations

  • Tevatron at LN2 temperature and LN2 usage is stable with vendor deliveries running only 2-3 truckloads per day due to the Tevatron being 1/4 at room temperature. Helium losses are steady at 15 kscf/day (1/2 of normal during TeV operations), some of the loss being replenished with A0 GHe Recovery from the test caves.
  • A1, A3, and D1 are at room temperature and isolated with LOTO. D1 has been certified by MSD and hipotted successfully thus Cryo will restore bldg soon.

B4, E2 and F4 are also at room temperature, have been reconnected by Cryo, and are under helium flow though Division Head permission has been granted for cooldown. B4 and E2 cooldown will be initiated after this weekend’s ACNET outage and Monday’s E-sector power outage. F4 cooldown will wait until TEL-1 is re-installed.

A3 has been completed by MSD, though spool piece conning towers may still be retrofitted (5 out of 8 leak) by Tech Division during this shutdown. Cryo has completed epoxy of helium leaks on Tevatron power leads, though they still need a snoop leak test.

  • Pbar D60 Debuncher vacuum work has been completed by Pbar Dept, thus Cryo personnel will restore helium/nitrogen connections and cool down tanks to LN2 temperature after Division sign-off.
  • ACNET application programs relating to FRIG operations continue to be tested on Linux consoles.
  • Nitrogen Reliquefier startup is delayed due to concern of “old” electrical cables being used for cross-connect between MSS and KRSS, especially due to high startup currents of 3000 hp TSI compressor. IR Booster compressor (400 hp) is running and purifying the plant.
  • MCR monitoring cryo operations.

CHL Maintenance:

  • General facility maintenance is in progress, including testing of relief valves per FESHM.
  • CHL Helium Compr “A” cooling tower installation has been completed with only some minor work is left, such as fencing; system testing will be commencing.
  • CHL Helium Compr “D” is being overhauled.
  • Helium Coldbox-2 controls upgrade is in progress.
  • Mobile Helium Purifier is being derimed at room temperature and piping modified.

FRIG Maintenance:

  • Various FRIG bldg vacuum maintenance and instrumentation repairs (e.g. lost charge on VPTs) are in progress, as well as support work on houses at room temperature and other equipment as things fail even during the shutdown period.
  • Various Mycom compressor maintenance: leak hunting and general repairs, F0#3 motor starter replacement in progress.
  • Wet expanders overhauls have been completed for A1 through F4; one D-Zero Dry Expander was also completed with modifications. AP30 and CDF remain to be done.
  • Kautzky relief valves have been replaced from C1 through F4 (16/24). Pace is at one house per day though running into problems with galling nuts on clamp's T-bolts and bad compression fittings. More Clamp T-bolts are on order.
  • We have been finding some leaking quick-disconnect fittings on conning towers. We have none in stock and thus order is being placed to get spares

Facility Engineering Services

  • CUB LCW pumps tested.
  • Zebra mussel cleaning, though waiting for a pproval, is scheduled to begin on May 8 and continue for a minimum of 16 days.
  • Casey’s Pond power failure test on Tuesday, 4-25-06.

Instrumentation

Work continues.

RF

Work continues.

Controls

Work continues.

ES&H Upgrades

Safety System Upgrades

Keep discussing safety with your crews to keep our focus on getting our jobs done safely.

  • Booster 60% Complete
    • Booster Dump Move Configuration Changes
      • MP01 and VBC1 Critical device control removed
      • LG1 area work next
    • Completing Documentation to submit for Approval to Operate
  • Main Injector 40% Complete
    • Lower Level Controlled Access Displays Replaced
      • MI-60 area complete. 7 Buildings left.
    • Main Injector Coasting Beam System changes
      • Lam10 Critical Device Interface Installation
    • Stairwell cleaning continues
      • MI-60 through MI-30 complete working toward MI-20.
  • P-Bar Source 100% Complete
    • Safety System Power Supply & Junction Box Replacements
    • Interlock New BSC700 Beam Stop
  • Tevatron 95% Complete
    • Moving location of both Tevatron Coasting Beam Gate Valves
    • F-Sector Radiation Monitor Card Replacement
  • MT6 Section 1 & 2 70% Complete
    • Enclosure Interlock System Upgrade
  • Safety System Testing 0% Complete
    • All areas need to be re-tested before startup
    • Booster testing is expected ~ May 1st
  • Continue to focus on getting shutdown work done safely.

EE Support

A0 kicker modifications continue.

Alignment

  • Tevatron
    • Scheduled As found two sectors.
  • Main Injector
    • Scheduled Wide Aperture Quad
  • Recycler
    • Scheduled Cooling Kicker Tank
    • Scheduled Damper Kicker
    • Scheduled Alignment suspect magnets
  • 8Gev
    • Scheduled To Complete Pipe and New Toroid
  • P-Bar
    • Schedule Girder Check
    • Scheduled ceramic beam tube alignment
  • MiniBooNE
    • Scheduled May 1
  • NuMI
    • Schedule Re-Survey Multiwire and OTR.
  • Booster
    • Not Scheduled
  • 400Mev
    • Not Scheduled
  • Add Ons
    • D0 - Make Level-Run vertical measurement of select Collision Hall monuments West Wall
    • Main Injector - Align barrier cavity at MI604 location.

CUB Outage Schedule

  • April 21, 2006
    • LINAC Shutdown For Outage
    • Computer Room Is Already Running On Temporary Chiller
  • April 22
    • 0600 CUB Water Turned Off
    • 0700 Feeder 42 (CUB Power Turned Off)
    • 0700 Feeder 40 Linac & XGallery Off (9hrs)
    • Computer Room Chiller Off, Computers?
    • 0701 CUB Generator On, Light & Sump power for Booster Tower West.
    • Directional Boring Starts, Update at 1400.
    • 1600 Feeder 40 work complete, Feeder On
    • 1630 Computer Room Chiller Restarted
    • 1630 Starting Up Computers
  • April 23
    • 0900 Feeder 42 Cable Hi-pot
    • 1200 Feeder 42 (CUB Power Turned On)
    • 1300 CUB Water Back Up & Ready
    • Computer Room Can Be Moved Off The Temporary Chiller To CUB.
  • Additional Information
    • There is a possibility that CUB may come up on Saturday Evening if things go really well.
    • There is a possibility that CUB may not be back Monday if a Feeder 42 cable is hit.
    • There is NO Master Substation Outage on Monday, April 24 th @ 0700 as previously planned.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/21/2006    email Fermilab